Compute the Wronskian of yı(t)= e? and yz(t)=e-37 -5e- -e- -5e- -61 -e-61
Answer #1

Option a is true

Q: Wlaonskian of y.lt)= 02 is A -34 Y 2 = e 1-oi-e 27 e e -3t c 2e24 -3t - 3e Yg (t) Yect) 4+1+) yy (t) = -3624-36 – 2824-34

Please rate this answer Thank you ?